how do u graph y=2x+5? not sure how 2 begin
Answer by wuwei96815(245)

You only need two coordinates to graph a linear equation because the graph is a straight line.

y=2x+5
Let x=0
y=2(0)+5=5
So, when x=0, y=5

y=2x+5
Let y=0
0=2x+5
-5=2x
-5/2=x
So, when y=0, x=-5/2=-2.5

You now have your two coordinates (0,5) and (-2.5,0). Plot the two points on a graph, draw a line through them, and you have your graph.